OpenStudy (anonymous):

How to solve for x in this equation e ^{-x} = 3e ^{-3x}

OpenStudy (anonymous):

\[e ^{-x} = 3e ^{-3x}\]

OpenStudy (inkyvoyd):

Take the natural log of both sides.

OpenStudy (anonymous):

The original function is e ^{-x} = e ^{-3x} and i did the first derivative and got the answer above

OpenStudy (inkyvoyd):

No, don't take the derivative.

OpenStudy (anonymous):

I got x= 0.55 but my book's answer is 0.38 or so i dont know if i have differnciate wrong or what

OpenStudy (anonymous):

the question asks for extreme values, so need to find first derivative

OpenStudy (inkyvoyd):

Oh, Nevermind.

OpenStudy (inkyvoyd):

You took the derivative incorrectly

OpenStudy (inkyvoyd):

should be -e^(-x)=-3e^(-3x)

OpenStudy (anonymous):

yea, that's what i have too and then set f'(x) to 0

OpenStudy (anonymous):

the 2 negative cancels out

OpenStudy (inkyvoyd):

Can you type out the exact question?

OpenStudy (anonymous):

Use algorithm for finding extreme values to determine the absolute maximum and minimum values of the function f(x) = \[e ^{-x} - x ^{-3x}\]

OpenStudy (anonymous):

sorry, that's e, not x

OpenStudy (inkyvoyd):

I'm getting a math processing error, leme swtich browsers.

OpenStudy (anonymous):

\[e ^{-x} - e ^{-3x}\]

OpenStudy (anonymous):

ok! thanks

OpenStudy (inkyvoyd):

y=e^(-x)-e^(-3x) y'=-e^(-x)+3e^(-3x) Since y'=0 e^(-x)=3e^(-3x) I got 0.55 too (wolfram alpha input) I think your textbook answer is wrong.

OpenStudy (anonymous):

ohhh i know why now because that 0.38 is when u subsitute 0.55 in the f (x) function! thanks for ur help!

OpenStudy (inkyvoyd):

0.o. I did absolutely nothing though.

OpenStudy (anonymous):

thankful for trying to help :)

OpenStudy (inkyvoyd):

oh :D